a) The pOH of an aqueous solution at 25°C was found to be 3.50.

The pH of this solution is .

The hydronium ion concentration is ___M.

The hydroxide ion concentration is ____M.

b) The pOH of an aqueous solution at 25°C was found to be 5.60.

The pH of this solution is .

The hydronium ion concentration is ___M.

The hydroxide ion concentration is ___M.

i'll give a good rating!
Answer #1

(at 25°c) sans cal given: pon = 3.50 we know рои = 14 ри. + So ри ри = 14- рой ри = 19 – 3-50 = To-So and we know ри. -log[и](6) given: pon = 5.60 5-60 (at 25 c) pH + pol = 14 pH = 14-pon pH = 14-5.6 pH = 8.4 So Hydronium ion concentradion is pH = -l